# RecJPQ: Training Large-Catalogue Sequential Recommenders

### 11 Dec 2023 v1

@@ 5.2. Results
Table 4 shows the experimental results on the smaller ML-1M and Booking datasets, while Table 5 reports results for the Gowalla dataset. The tables compare NDCG@10 and model size of compressed variations of backbone models with the base (uncompressed) model. Significant differences compared to the corresponding base model (BERT4Rec, GRU or SASRec) are indicated. In general, the tables show that RecJPQ substantially reduces the model checkpoint size in most cases. For example, the RecJPQ versions of the GRU models on the Booking dataset are approximately 27% of the original in size. On the Gowalla dataset, compressed models are approximately 3% of the original. Moreover, model size does not depend on the centroid assignment strategy. Indeed, centroid assignments only influence the values of the model parameters but not the number of parameters. Moreover, Quotient Remainder models have approximately the same compression level as RecJPQ models. We speculate that after compression, the model checkpoint size is dominated by other model parameters (e.g., attention matrices). In our configuration, the centroid embeddings tensor only requires a few megabytes of memory (see Table 2). In contrast, the full model checkpoint of a compressed model is typically tens of megabytes (e.g., 92.8MB for SASRec using RecJPQ-BPR trained on Gowalla).
@@
RecJPQ only increased the model size for SASRec on MovieLens-1M due to the dataset’s small item count. The overhead of storing centroid embeddings and the codebook eclipses the benefit of compressing the embeddings table. Using RecJPQ with smaller embeddings might reduce the model size without affecting performance on this dataset (see Section 5.2.3).
202020
@@
On the other hand, we observe from Table 4 and Table 5 that the choice of the optimal strategy depends on both the model and the dataset. For example, on MovieLens-1M, the choice of the strategy is not important, and in all cases, RecJPQ versions of the models are statistically indistinguishable from the base for all models. On the larger Booking dataset, the choice of the best strategy is model-dependent. For BERT4Rec, the best results are achieved with BPR (NDCG@10 0.375, statistically indistinguishable from the base) and SVD (NDCG@10 0.379, +0.97%, significant). At the same time, the Random strategy significantly underperforms the base configuration (NDCG@10 0.316, -15.98%) - this shows that in some cases, assigning similar codes to similar items is indeed important. However, in 2 cases, Random performs statistically significantly better than SVD and BPR. For example, Random assignments perform best on Gowalla with the SASRec base (a significant improvement of +57% over the base). SVD assignments also moderately improve the result in this case (+10%, significant). At the same time, BPR decreases the quality by a large margin on Gowalla dataset (-70%) [^21] The percentages for the Gowalla dataset seem large because this dataset is difficult: it has the largest number of items and the largest proportion of long-tail items.. We explain the success of the Random strategy on the Gowalla dataset as giving a larger regularisation effect (random assignments make the learning task harder, so the model has fewer chances to overfit). Overall, from Tables 4 and 5, we see that there is no ”one size fits all” choice of the centroid assignment strategy, and it depends on both dataset and model salient characteristics; the exact best combination of model/strategy may depend on regularisation requirements (as we observe in Gowalla), the prevalence of strong sequential patterns (as in Booking, see (Petrov and Macdonald, 2022a) for details) and so on. This suggests that the centroid assignment strategy could be treated as a hyperparameter and tuned for each model/dataset combination. However, by default, we recommend using RecJPQ with SVD strategy - in all cases, it achieves significantly better (on the Booking and Gowalla datasets) or statistically indistinguishable (on the MovieLens-1M dataset) results compared to the base model. We also note that RecJPQ with the SVD strategy is always better than the Quotient Remainder baseline (Quotient Reminder is always significantly worse than the base mode, whereas RecJPQ is better or indistinguishable). The question of whether or not it is possible to select the best model/strategy combination without doing an exhaustive hyperparameter search is an interesting research direction, which we leave for future work.
@@
Table 5. Impact of RecJPQ with different centroid assignment strategies on SASRec model size and effectiveness on the large-scale Gowalla dataset. Notations follow Table 4.
@@
It is also worth mentioning that we do not observe a degradation of the training efficiency when training the RecJPQ versions of the models. Indeed, while there are some fluctuations in the training time the model requires to converge, the magnitude of the required training time remains the same: for example, training of the base version of BERT4Rec requires 18.8 hours on Booking.com, and the RecJPQ-SVD version of BERT4Rec requires 16.1 hours. The training time of the SVD model (used for initial centroid assignments) in the same case is negligible compared to the training of the main model (it takes approximately a minute). Inference time is also unaffected (e.g. on Gowalla, full evaluation across the 86k users requires 10 minutes for both ”base” and RecJPQ versions of the models).
In summary, in answer to RQ1, we conclude that RecJPQ achieves large model compression levels. The achieved compression is particularly impressive on datasets with large catalogues (like Gowalla). The compression does not depend on the centroid assignment strategy. However, the centroid assignment strategy greatly affects the model performance. The effect is model- and dataset-dependent, so the strategy should be treated as a hyperparameter. However, the SVD is a safe choice, as it always provides results that are comparable (i.e statistically indistinguishable) or better than the base model.
@@
Figure 3. RecJPQ performance while varying embedding size $d$ and the number of centroids per item $m$.
@@
To answer RQ2, we perform a grid search over embedding size and code length on the MovieLens-1M and Gowalla datasets. We use SASRec as the backbone (the only model that can be easily trained on Gowalla) and apply the SVD centroid assignment strategy. We select the embedding size $d$ from {$2^{0},2^{1},2^{2},...,2^{9}$} and code length $m$ from {$2^{0},2^{1},2^{2},...,2^{8}$}. Note that $m\leq d$, as RecJPQ splits each embedding of size $d$ into $m$ sub-embeddings.
222222
@@
Figure 3 illustrates the results of the grid search. The figure shows the NDCG@10 of the SASRec-RecJPQ model for each combination of code length (x-axis) and embedding size (y-axis), in the form of a heatmap for both datasets. As we can see from the figure, a larger embedding size generally positively affects the model performance. This result echoes similar findings of a recent reproducibility paper (Rendle et al., 2022); however, interestingly, in the RecJPQ case, increasing embedding dimensionality does not change the amount of information we store per each item, as the length of the code defines it rather than the embedding size. Instead, it increases model capacity increasing the amount of information that can be stored in each centroid, allowing centroids to account for more item characteristics. For example, on Gowalla, the largest embedding we can train using base SASRec is 128-dimensional embedding, while with RecJPQ, we can train the model even with 512-dimensional embeddings.
@@
On the other hand, larger code lengths is not always helpful. As we discussed in Section 4.2, RecJPQ forces the model to share parts of embeddings with other items and therefore acts as a regularisation mechanism. Shorter code length forces items to share more information, and therefore it causes a stronger regularisation effect. As we can see, on the less sparse MovieLens-1M – where all items have more than five interactions – regularisation is not an issue, and longer codes are beneficial. For example, the best result is achieved with 512-dimensional embeddings and a code of length 128 (NDCG@10 0.14). In contrast, for Gowalla, where most items are long-tail items with less than five interactions (hence the embeddings of these items should be regularised), the best NDCG is achieved with the code of length 8 (NDCG@10 0.12). The fact that the model can perform better with shorter codes confirms that RecJPQ can work as a regularisation technique.
@@
Figure 4. Model Performance/Model Size tradeoff for SASRec and SASRec-RecJPQ.
To address our last research question, we analyse the trade-off between model checkpoint size and NDCG@10 achieved by the model when trained with different embedding sizes. We select the embedding size from {1, 2, 4, 8, 16, 32, 64, 128, 256, 512} and train the original versions of SASRec and SASRec-RecJPQ with the SVD strategy on the MovieLens-1M and Gowalla datasets. For RecJPQ, we select code length $m$ optimal for the dataset/embedding size pair (according to grid search from Section 5.2.2). For the original SASRec on Gowalla, we only train up to the embedding size of 128 (larger embedding sizes cannot be trained on our GPUs’ memory).
232323
@@
Overall in answer to RQ3, we conclude that while larger models benefit model performance, RecJPQ improves this tradeoff by a large margin (i.e. to achieve the same performance, RecJPQ requires much fewer parameters than the original model). This effect is more markedly pronounced for the larger Gowalla dataset.
@@ 6. Conclusions
In this paper, we discussed the challenge of training sequential recommender systems with large datasets, primarily due to the large item embedding tensor. Existing compression methods have limitations, leading to the proposed method, RecJPQ, based on Joint Product Quantisation. Our evaluation of RecJPQ on three datasets resulted in significant model size reduction, e.g., 47.94$\times$ compression of the SASRec model on the Gowalla dataset. Additionally, RecJPQ serves as a model regulariser, improving model quality, with SASRec-RecJPQ using SVD strategy outperforming the original SASRec model (+35% NDCG@10 on Booking, +10% on Gowalla). This paper’s method could bridge the gap between academic research, especially on transformer-based architectures, and large-scale production recommender systems of large companies. There are also a number of future research directions, which can be based on this work, including deep analysis of RecJPQ on training/inference time and analysing theoretical grounds for the observed regularisation effects.

### 18 Dec 2023 v2

@@ 5.2. Results
RQ1. Effect of sub-item id assignment strategy. Table 4 shows the experimental results on the smaller ML-1M and Booking datasets, while Table 5 reports results for the Gowalla dataset. The tables compare NDCG@10 and model size of compressed variations of backbone models with the base (uncompressed) model. Significant differences compared to the corresponding base model (BERT4Rec, GRU or SASRec) are indicated. In general, the tables show that RecJPQ substantially reduces the model checkpoint size in most cases. For example, the RecJPQ versions of the GRU models on the Booking dataset are approximately 27% of the original in size. On the Gowalla dataset, compressed models are approximately 3% of the original. Moreover, model size does not depend on the sub-item id assignment strategy. Indeed, sub-item id assignments only influence the values of the model parameters but not the number of parameters. Moreover, Quotient Remainder models have approximately the same compression level as RecJPQ models. We speculate that after compression, the model checkpoint size is dominated by other model parameters (e.g., attention matrices). In our configuration, the sub-item embeddings tensor only requires a few megabytes of memory (see Table 2). In contrast, the full model checkpoint of a compressed model is typically tens of megabytes (e.g., 92.8MB for SASRec using RecJPQ-BPR trained on Gowalla).
161616
@@
RecJPQ only increased the model size for SASRec on MovieLens-1M, due to the dataset’s small item count: on this dataset, the overhead of storing sub-item embeddings and the codebook is larger compared to the benefit of compressing the embeddings table. Using RecJPQ with smaller embeddings might reduce the model size without affecting effectiveness on this dataset (see also RQ3).
@@
On the other hand, we observe from Table 4 and Table 5 that the choice of the best assignment strategy depends on both the model and the dataset. For example, on MovieLens-1M, the choice of the strategy is not important, and in all cases, RecJPQ versions of the models are statistically indistinguishable from all corresponding base models. On the larger Booking dataset, the choice of the best strategy is model-dependent. For BERT4Rec, the best results are achieved with BPR (NDCG@10 0.375, statistically indistinguishable from the base) and SVD (NDCG@10 0.379, +0.97%, significant). At the same time, the Random strategy significantly underperforms the base configuration (NDCG@10 0.316, -15.98%) – this shows that in some cases, assigning similar codes to similar items is indeed important. However, in 2 cases, Random performs statistically significantly better than SVD and BPR. For example, Random assignments perform best on Gowalla with the SASRec base (a significant improvement of +57% over the base). SVD assignments also moderately improve the result in this case (+10%, significant). At the same time, BPR decreases the quality by a large margin on Gowalla dataset (-70%) [^17] The percentages for the Gowalla dataset seem large because this dataset is difficult: it has the largest number of items and the largest proportion of long-tail items.. We explain the success of the Random strategy on the Gowalla dataset as giving a larger regularisation effect (random assignments make the learning task harder, so the model has fewer chances to overfit). Overall, from Tables 4 and 5, we conclude that there is no ”one size fits all” choice of the sub-item id assignment strategy, and it depends on both dataset and model salient characteristics; the exact best combination of model/strategy may depend on regularisation requirements (as we observe in Gowalla), the prevalence of strong sequential patterns (as in Booking, see (Petrov and Macdonald, 2022a) for details) and so on. This suggests that the sub-item id assignment strategy could be treated as a hyperparameter and tuned for each model/dataset combination. However, by default, we recommend using RecJPQ with the SVD strategy – in all cases, this achieves significantly better (on the Booking and Gowalla datasets) or statistically indistinguishable (on the MovieLens-1M dataset) results compared to the base model. We also note that RecJPQ with the SVD strategy is always better than the Quotient Remainder baseline (Quotient Reminder is always significantly worse than the base model, whereas RecJPQ is better or indistinguishable). The question of whether or not it is possible to select the best model/strategy combination without doing an exhaustive hyperparameter search is an interesting research direction, which we leave for future work.
@@
Table 5. Impact of RecJPQ with different sub-item id assignment strategies on SASRec model size and effectiveness on the large-scale Gowalla dataset. Notations follow Table 4.
@@
Figure 3. RecJPQ performance for SASRec while varying embedding size $d$ and the number of sub-item ids per item $m$.
181818
It is also worth mentioning that we do not observe a degradation of the training efficiency when training the RecJPQ versions of the models. Indeed, while there are some fluctuations in the training time the model requires to converge, the magnitude of the required training time remains the same: for example, training of the base version of BERT4Rec requires 18.8 hours on Booking.com, and the RecJPQ-SVD version of BERT4Rec requires 16.1 hours. The training time of the SVD model (used for initial sub-item id assignments) in the same case is negligible compared to the training of the main model (it takes approximately a minute). Inference time is also unaffected (e.g. on Gowalla, full evaluation across the 86k users requires 10 minutes for both ”base” and RecJPQ versions of the models).
In summary, in answer to RQ1, we conclude that RecJPQ achieves large model compression levels. The compression is particularly impressive on datasets with large catalogues (like Gowalla). The compression does not depend on the sub-item id assignment strategy. However, the sub-item id assignment strategy greatly affects the model performance. The effect is model- and dataset-dependent, so the strategy should be treated as a hyperparameter. However, the SVD is a safe choice, as it always provides results that are comparable (i.e statistically indistinguishable) or better than the base model.
@@
RQ2. Effects of code length $m$ and the embedding size on model performance. To answer RQ2, we perform a grid search over embedding size and code length on the MovieLens-1M and Gowalla datasets. We use SASRec as the backbone (the only model that can be easily trained on Gowalla) and apply the SVD sub-item id assignment strategy. We select the embedding size $d$ from {$2^{0},2^{1},...,2^{9}$} and code length $m$ from {$2^{0},2^{1},...,2^{8}$}. Note that $m\leq d$, as RecJPQ splits each embedding of size $d$ into $m$ sub-embeddings.
191919
@@
Figure 3 illustrates the results of the grid search. The figure shows the NDCG@10 of the SASRec-RecJPQ model for each combination of code length (x-axis) and embedding size (y-axis), in the form of a heatmap for both datasets. As we can see from the figure, a larger embedding size generally positively affects the model performance. This result echoes similar findings of a recent reproducibility paper (Rendle et al., 2022); however, interestingly, in the RecJPQ case, increasing embedding dimensionality does not change the amount of information we store per each item, as the length of the code defines it rather than the embedding size. Instead, it increases model capacity, increasing the amount of information that can be stored in each sub-item embedding, allowing to account for more item characteristics. For example, on Gowalla, the largest embedding we can train using base SASRec is 128 dimensions, while with RecJPQ, we can train the model even with 512-dimensional embeddings.
@@
On the other hand, larger code lengths are not always helpful. As we discussed in Section 4.2, RecJPQ forces the model to share parts of embeddings with other items and, therefore, acts as a regularisation mechanism. A shorter code length forces items to share more information, and therefore, it causes a stronger regularisation effect. As we can see, on the less sparse MovieLens-1M – where all items have more than five interactions – regularisation is not an issue, and longer codes are beneficial. For example, the best result is achieved with 512-dimensional embeddings and a code of length 128 (NDCG@10 0.14). In contrast, for Gowalla, where most items are long-tail items with less than five interactions (hence the embeddings of these items should be regularised), the best NDCG is achieved with the code of length 8 (NDCG@10 0.12). The fact that the model can perform better with shorter codes confirms that RecJPQ can behave as a regularisation technique.
202020
@@
Figure 4. NCDG/Size tradeoff for SASRec and SASRec-RecJPQ.
RQ3. Size-Performance tradeoff. To address our last research question, we analyse the trade-off between model checkpoint size and NDCG@10 achieved by the model when trained with different embedding sizes. We select the embedding size from {1, 2, 4, 8, 16, 32, 64, 128, 256, 512} and train the original versions of SASRec and SASRec-RecJPQ with the SVD strategy on the MovieLens-1M and Gowalla datasets. For RecJPQ, we select code length $m$ that is optimal for the dataset/embedding size pair (according to the grid search from Figure 3). For the original SASRec on Gowalla, we only train up to the embedding size of 128 due to GPU memory limit.
@@
Overall, in answer to RQ3, we conclude that while larger models benefit model performance, RecJPQ improves this tradeoff by a large margin (i.e. to achieve the same performance, RecJPQ requires much fewer parameters than the original model). This effect is more markedly pronounced for the larger Gowalla dataset.
@@ 6. Conclusions
In this paper, we discussed the challenge of training sequential recommender systems with large datasets, primarily due to the large item embedding tensor. Existing embedding compression methods have limitations, leading to our proposed method, RecJPQ. Our evaluation of RecJPQ on three datasets resulted in significant model size reduction, e.g., 47.94$\times$ compression of the SASRec model on the Gowalla dataset. Additionally, RecJPQ serves as a model regulariser, improving the model’s quality, with SASRec-RecJPQ using SVD strategy outperforming the original SASRec model (+35% NDCG@10 on Booking, +10% on Gowalla). There are a number of future work directions, which can be based on this paper, e.g., an analysis of the RecJPQ effect on training/inference time and analysing the theoretical grounds for the observed regularisation effects.